Zink has a work function of 4.3 eV . What is the longest wavelength of light that will release an electron from a zink surface? A 4.6 eV photon strikes the surface of zink and an electron is emitted. What is the maximum possible speed of the electron?
Answer #1

SOLUTION :

===> We have

E =1243eV/λnm

W =4.3 eV

Longest wavelength of light

λ = 1243/4.3

= 290 nm

===>

E =4.6 eV

Mass of electron

m =9.1x10^-31 kg

e =1.6x10^-19C

From Einstein photo electric effect equation

E =W+Kmax

Kmax = (1/2)mvmax^2 = E-W

(1/2)mvmax^2 = 4.6 -4.3 =0.3 eV

vmax^2 = (2x0.3x1.6x10-19)/(9.1x10-31) =0.105 x1012

vmax = 3.24 x105 m/s